OpenStudy (anonymous):

Solve the following. 3.0 (x - 2.5) = 7.5

OpenStudy (anonymous):

(5q - 35m + 30q) - (11n + m + 7q) (35q - 35m) - (11n + m + 7q) 28q - 36m - 11n

OpenStudy (anonymous):

so you dont multiply the 5 by the second brackets right? I forgot how to do this :(

OpenStudy (anonymous):

right, it is only part of the first series of variables, if all of them were in one set of parenthesis then you would multiply all by 5

OpenStudy (anonymous):

alright thank you :) do you know this by the way? Solve the following. a) 3.0 (c-2.5) = 7.5

OpenStudy (anonymous):

no prob and lets see.. c = 5

OpenStudy (anonymous):

oh shoot sorry, it's not a "c" its an "x"

OpenStudy (anonymous):

oh okay well same thing haha

OpenStudy (anonymous):

Alright, sorry

OpenStudy (anonymous):

Would it be.... 3.0(x-2.5)=7.5 3.0x - 7.5 = 7.5 ?

OpenStudy (anonymous):